# NeuroBench: A Framework for Benchmarking Neuromorphic Computing Algorithms and Systems

### v3

@@ Acknowledgements
This work has been supported in parts by Semiconductor Research Corporation (SRC), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001448), a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No. CityU 11200922], ARC Laureate Fellowship FL210100156, and the EU H2020 project BeFerroSynaptic (871737). The authors would like to acknowledge the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).
@@
Authors are grouped based on contributions, and ordered alphabetically within groups. JY led project discussions and management. JY and KVdB implemented harness metric infrastructure, conducted experiments, and prepared the manuscript. The following authors primarily developed the main results: DdB and MF on the keyword few-shot class-incremental learning task; GT and SW on the event camera object detection task; PH, PVS and BZ on the non-human primate motor prediction task; YB and DK on the chaotic function prediction task; and NP led harness infrastructure development. SHA, GVJ, BL, AM, AKM, GL, and TS developed components of the harness infrastructure. ZA, MA, BA, AGA, CB, AB, PB, SB, SB, GC, EC, FC, GdC, AD, AD, MD, YD, JE, TF, JF, VF, SF, PMF, WG, AG, HAG, GI, SJ, VK, LK, JCK, LK, RK, DK, YL, SL, HM, RM, FMM, CM, KM, DM, EN, TN, FO, AO, PP, JP, MP, CP, MAP, AP, CP, AR, YS, CJSS, AvS, JS, SS, CS, JS, SS, SBS, MS, AS, MS, KS, TCS, PS, JT, NT, GU, MV, CMV, BV, AY, and FTZ participated in discussions during meetings, prepared sections for the present manuscript and/or its preprint, and reviewed the manuscript. CF and VJR jointly supervised the project, reviewed the manuscript, and analyzed results.

### v5

@@ Acknowledgements
Authors of this work have been supported in parts by Semiconductor Research Corporation (JY), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001448), a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No. CityU 11200922], ARC Laureate Fellowship FL210100156, and the EU H2020 project BeFerroSynaptic (871737). We acknowledge the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen). We acknowledge a contribution from the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union –NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”). The work of SynSense was partially supported by the European Commission, under the Horizon grant Ferro4Edge AI (grant agreement 101135656). This work is partly funded by the German Federal Ministry of Education and Research (BMBF) and the free state of Saxony within the ScaDS.AI center of excellence for AI research and by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) under contract 01MN23004F (ESCADE).
@@
Authors are grouped based on contributions, and ordered alphabetically within groups. JY led project discussions and management. JY and KVdB implemented harness metric infrastructure, conducted experiments, and prepared the manuscript. The following authors primarily developed the main algorithm track results: DdB and MF on the keyword few-shot class-incremental learning task; GT and SW on the event camera object detection task; PH, PVS and BZ on the non-human primate motor prediction task; YB and DK on the chaotic function prediction task; and NP led harness infrastructure development. The following authors primarily developed the main system track results: WK and MAK on the acoustic scene classification task; AP and PS on the QUBO task. SHA, GVJ, BL, AM, AKM, GL, and TS developed components of the algorithm track harness infrastructure. ZA, MA, BA, AGA, CB, AB, PB, S. Bohte, S. Buckley, GC, EC, FC, GdC, A. Danielescu, A. Daram, MD, YD, JE, TF, JF, VF, SF, PMF, WG, AG, HAG, GI, SJ, VK, L. Khacef, JCK, L. Kriener, RK, DK, SL, YL, HM, RM, JMM, CM, KM, DRM, EN, TN, FO, AO, PP, JP, MP, C. Pehle, MAP, C. Posch, AR, YS, CJSS, AvS, J. Schemmel, S. Schmidgall, CS, J. Seo, S. Sheik, SBS, MS, AS, KS, MS, TCS, JT, NT, GU, MV, CMV, BV, AY, and FTZ participated in discussions during meetings, prepared sections for the present manuscript and/or its preprint, and reviewed the manuscript. CF and VJR jointly supervised the project, reviewed the manuscript, and analyzed results.